OpenStudy (anonymous):

Simplify. Assume all variables are possible 3sqrt-54x^6y^5

OpenStudy (anonymous):

is this the question- \(\Large \sqrt[3]{54x^6y^5}\)

OpenStudy (anonymous):

yes but the 54 is negative!

OpenStudy (anonymous):

@hannahleigh633

OpenStudy (anonymous):

@at0m the 54 is negative

OpenStudy (anonymous):

oh okay so its this-> \(\Large \sqrt[3]{-54x^6y^5}\) right? :)

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

we can write it like this-> \(\Large \sqrt[3]{-2 \times 3^3} \times \sqrt[3]{x^6} \times \sqrt[3]{y^3 \times y^2}\) => \(\Large 3 \sqrt[3]{-2} \times x^2 \times y \sqrt[3]{y^2}\) => \(\Large 3x^2y \sqrt[3]{-2y^2}\)

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

Can you help with another? 3sqrt-192x^6y^5